# Observational Scaling Laws and the Predictability of Language Model Performance

### v1

@@ 1 Introduction
Although the high costs of compute scaling laws are unavoidable when optimizing pre-training hyperparameters (e.g., Hoffmann et al. [35]), this is not true of all scaling analyses. In this work, we show that many other types of scaling studies, such as understanding complex model capabilities (e.g. agentic or “emergent” behaviors) and post-training interventions, can be done using a lower-cost, higher-resolution, and broader-coverage alternative to the standard approach of training (or using) a single family of LMs across compute scales.
Figure 1: Observational scaling laws generalize existing compute scaling laws which directly relate training compute to downstream capabilities (dashed line) by hypothesizing the existence of a low-rank space of LM capabilities that have a log-linear relationship with compute (center), and can be predicted directly from standardized LM benchmarks (left). This enables us to get low-cost, high-resolution scaling predictions using nearly 80 publicly available LMs (right).
@@
Through an analysis of existing standardized LM benchmarks (e.g., Open LLM Leaderboard [9]), we find a few such capability measures that have scaling law relationships with compute within model families ($R^{2}>0.9$) (Fig. 3), and with downstream metrics across model families. We call such scaling relationships observational scaling laws as they relate simple observable quantities that we expect to scale with compute (such as standardized benchmark performance) with complex downstream quantities of interest.
The ability to build scaling laws across a large number of existing models has significant advantages in cost, resolution, and coverage: observational scaling incurs no training cost, while leveraging a large number of models spanning a much larger compute range than any single model family. Observational scaling also significantly increases the resolution of scaling laws by virtue of using more models, which is useful for studying nearly discontinuous phenomena like “emergent” capabilities. Finally, observational scaling can combine model families from heterogeneous sources with very different scaling properties and capabilities (e.g., LLaMA [83] vs StarCoder [46]) which allows us to study how different scaling strategies impact downstream performance and algorithmic interventions.
@@
Emergent capabilities (Sec. 4.1)  There has been an active debate about whether LMs have “emergent” capabilities that discontinuously appear at certain compute thresholds and whether these capabilities can be predicted using small models [90, 74, 56, 36]. The high resolution of observational scaling laws show that some of these phenomena follow a smooth sigmoid, and can be predicted accurately using small, sub Llama-2 7B models.
Agentic capabilities (Sec. 4.2)  We show that the more high-level, complex capabilities of LMs as agents, as measured by AgentBench [54] and AgentBoard [57], can be predicted using observational scaling laws. Our scaling law precisely predicts the performance of GPT-4 using only weaker models (sub GPT-3.5) and identifies programming capabilities as driving agent performance.
Post-training method scaling (Sec. 4.3)  We show that our scaling laws can reliably predict the gains of post-training techniques, such as Chain-of-Thought [91] and Self-Consistency [89] at scale, even when we fit our scaling laws on weak models (sub Llama-2 7B).
@@ 2 Related Work
Scaling laws have been generalized beyond pretraining loss to analyze transfer learning [33, 78, 1] and downstream performance [32, 29, 15] across various domains, see Villalobos [86] for a comprehensive review. In particular, there has been evidence suggesting that the few-shot performance of LMs on downstream benchmarks is closely tied to compute measures like model size [13], but whether this is predictable with scaling laws remains debated. Extensive research has explored the difficulties of predicting benchmark performance due to their appearing rapid “emergence” [90, 76, 26], while recent works argued the discontinuity is due to the metrics used [74, 56] or the lack of data points [36] (see Anwar et al. [5] for a survey on this topic). Finnveden [24] and Owen [63] have investigated the use of linear and sigmoidal scaling laws, derived from pretraining loss or computational measures, to extrapolate the benchmark performance. Recent studies have also more extensively investigated the correlations between the pretraining loss and downstream performance of LMs [93, 37], aiding in the understanding of downstream scaling [25] and emergent capabilities [23] of LMs. On the theory front, Arora and Goyal [6] derived a theory characterizing how performance on complex skills of LMs can be derived as a composition of base skills. While our work shares similar goals in that we aim to understand the downstream, post-training performance of models, we differ in our approach in that we aim to build practical higher-resolution scaling laws using multiple model families and their observable standard benchmark metrics.
@@ 3.1 Generalizing Compute Scaling Laws
We now show that the extracted PC measures $S$ scale log-linearly with training FLOPs within each model family, and can thus be interpreted as a cross-model generalization of compute scale $C$.
@@
3.4 Fitting Observational Scaling Laws
@@
In the multi-model family case, variations in compute efficiency mean that FLOPs and capabilities are no longer log-linear across model families. However, we can map all of the models to a shared, FLOPs-based capability measure using a metric we call $f$-equivalent FLOPs. The core idea of the approach is to represent each model’s capabilities by the following hypothetical: “how many log-FLOPs ($\log(\bar{C}_{m,f})$) would it take for a model in a family $f$ to match a model $m$”. We call $\log(\bar{C}_{m,f})$ the $f$-equivalent FLOP for model $m$, as it represents the performance of model $m$ relative to models in the reference model family $f$. This measure can be computed fairly easily as
@@ 4 Validating Observational Scaling Laws
As baselines, we compare our scaling predictions to existing compute-based scale measures like training FLOPs and model size. We used the mean squared error (MSE) on the holdout set as our main evaluation measure, as the target range is always normalized (0 to 1), and estimating the marginal variance in $R^{2}$ can add additional noise when the test set sizes are small.
@@ 4.1 Predictability of “Emergent” Capabilities
Figure 4: “Emergent” capabilities of LMs can be accurately predicted from weaker models to stronger ones with observational scaling laws, and using PC measures provides much more accurate predictions than using compute scale measures like training FLOPs and model size (see Fig. C.10). Two non-arithmetic and two arithmetic tasks from BigBench [75], which are identified as “emergent” in [90], are used for illustration.
@@ 5 Selecting Low-Cost Model Subsets for Practical Scaling Analyses
8 models Llama-2 {7B, 13B, 70B}, Mixtral {8x7B}, XGLM {564M, 1.7B, 4.5B, 7.5B}
@@
Qwen {7B, 14B, 72B}, OPT {125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 30B, 66B},
GPT-J {6B}, GPT-Neo {125M, 1.3B, 2.7B, 20B}
@@ 6 Discussion and Other Applications of Observational Scaling
Many existing benchmarks suffer from a limited dynamic range: they either saturate quickly for large models (e.g., HellaSwag, Winogrande) or have completely random performance for small models (e.g., MMLU, GSM8K), see Fig. C.4 for the behavior of each benchmark. In contrast, we find that PC-1 is a smooth capability measure that can be used to compare LMs across many orders of magnitude (at least 10 nats). This allows us to compare models from heterogeneous sources and of extremely different capabilities on a single, unified scale (Fig. 9). We believe that the high dynamic range of PC1 may make it suitable as an optimization target for pretraining, where architecture or data interventions can be benchmarked against PC-1 at small scales and validated at large scales.
@@ C.2 Properties of PC measures
In Fig. 9, we have shown that PC-1 can serve as a smooth capability measure for LMs that provide meaningful readouts across many orders of scales (about 10 nats). In Fig. C.4, we show that using a single benchmark metric as LM capability measures amy suffer from a limited dynamic range. In particular, they may either saturate quickly for large models (e.g., HellaSwag, Winogrande) or provide random readouts for weak models (e.g., MMLU, GSM8K).
@@ C.3 Robustness Checks
Figure C.8: Comparing different scale measures under different holdout cutoffs on post-training analysis tasks in Sec. 4.3. The training/test data size is varied by changing the FLOPs cutoff and the area under the test error curves (AUE) is used to measure the overall prediction errors. PC measures (with # = 3) consistently lead to an earlier transition to low prediction error region and much lower AUE compared to compute measures like training FLOPs and model size.
@@ C.6 Model Subset Selection
(d) 16 models
Figure C.16: Prediction results with different numbers of randomly selected models. The prediction performance is much worse than our selection method, even when 16 models are being selected.
@@ C.7 Fited Functional Forms for Preregistration of Predictions
“Emergent” capabilities (Sec. 4.1) Word Unscramble $&\phi^{-1}(Y,1.00)\\ =~{}&0.87\log\left(\bar{C}_{\text{Llama-2}}\right)-6.12\\ =~{}&6.75\text{PC1}-3.21\text{PC2}-1.37\text{PC3}-4.93\\ =~{}&1.03\text{MMLU}+3.02\text{ARC-C}+5.73\text{HellaSwag}+2.44\text{Winogrande}~{}-\\ &1.06\text{TruthfulQA}+1.21\text{GSM8K}+2.48\text{XWinogrande}-0.08\text{HumanEval}-12.28\end{aligned}$
Persian QA $&\phi^{-1}(Y,1.00)\\ =~{}&1.02\log\left(\bar{C}_{\text{Llama-2}}\right)-8.46\\ =~{}&2.98\text{PC1}+3.37\text{PC2}+1.33\text{PC3}-5.29\\ =~{}&1.88\text{MMLU}+1.14\text{ARC-C}+0.91\text{HellaSwag}+0.43\text{Winogrande}~{}+\\ &0.32\text{TruthfulQA}+2.50\text{GSM8K}+1.22\text{XWinogrande}+2.88\text{HumanEval}-8.19\end{aligned}$
3-Digit Substraction $&\phi^{-1}(Y,1.00)\\ =~{}&2.39\log\left(\bar{C}_{\text{Llama-2}}\right)-8.92\\ =~{}&5.98\text{PC1}+8.74\text{PC2}+39.54\text{PC3}-4.68\\ =~{}&2.17\text{MMLU}+2.32\text{ARC-C}-3.44\text{HellaSwag}-7.96\text{Winogrande}~{}+\\ &0.65\text{TruthfulQA}+34.27\text{XWinogrande}+20.39\text{HumanEval}-20.99\end{aligned}$
2-Digit Multiplication $&\phi^{-1}(Y,1.00)\\ =~{}&0.96\log\left(\bar{C}_{\text{Llama-2}}\right)-4.45\\ =~{}&3.60\text{PC1}+4.24\text{PC2}+8.05\text{PC3}-2.68\\ =~{}&1.62\text{MMLU}+1.95\text{ARC-C}+0.55\text{HellaSwag}-0.63\text{Winogrande}~{}+\\ &0.14\text{TruthfulQA}+6.80\text{XWinogrande}+6.52\text{HumanEval}-8.00\end{aligned}$
Agentic capabilities (Sec. 4.2) AgentBench $&\phi^{-1}(Y,0.99)\\ =~{}&0.44\log\left(\bar{C}_{\text{Llama-2}}\right)-5.57\\ =~{}&2.34\text{PC1}+0.82\text{PC2}+2.94\text{PC3}-2.98\\ =~{}&2.46\text{MMLU}+0.85\text{ARC-C}+0.25\text{HellaSwag}+0.58\text{Winogrande}~{}+\\ &0.65\text{TruthfulQA}-0.48\text{GSM8K}+2.63\text{HumanEval}-6.45\end{aligned}$
AgentBoard $&\phi^{-1}(Y,0.97)\\ =~{}&0.43\log\left(\bar{C}_{\text{Llama-2}}\right)-6.60\\ =~{}&3.02\text{PC1}+2.60\text{PC2}-1.17\text{PC3}-2.98\\ =~{}&-0.10\text{MMLU}-0.31\text{ARC-C}-0.55\text{HellaSwag}+0.14\text{Winogrande}\\ &+0.56\text{TruthfulQA}+2.28\text{GSM8K}+3.36\text{HumanEval}-5.06\end{aligned}$
@@
Post-training (analysis Sec. 4.3) GSM Naive + Greedy $&\phi^{-1}(Y,1.00)\\ =~{}&0.49\log\left(\bar{C}_{\text{Llama-2}}\right)-4.88\\ =~{}&2.79\text{PC1}+1.66\text{PC2}+0.25\text{PC3}-3.62\\ =~{}&1.52\text{MMLU}+1.33\text{ARC-C}+1.28\text{HellaSwag}+0.76\text{Winogrande}~{}+\\ &0.13\text{TruthfulQA}+0.24\text{XWinogrande}+2.08\text{HumanEval}-6.16\end{aligned}$
GSM CoT + Greedy $&\phi^{-1}(Y,1.00)\\ =~{}&0.89\log\left(\bar{C}_{\text{Llama-2}}\right)-5.54\\ =~{}&2.56\text{PC1}+4.63\text{PC2}+4.30\text{PC3}-2.50\\ =~{}&5.08\text{MMLU}+2.04\text{ARC-C}-0.12\text{HellaSwag}+0.95\text{Winogrande}~{}+\\ &1.77\text{TruthfulQA}-2.42\text{XWinogrande}+2.55\text{HumanEval}-4.76\end{aligned}$
GSM CoT + SC $&\phi^{-1}(Y,1.00)\\ =~{}&0.95\log\left(\bar{C}_{\text{Llama-2}}\right)-5.75\\ =~{}&2.71\text{PC1}+4.81\text{PC2}+5.04\text{PC3}-2.49\\ =~{}&5.64\text{MMLU}+2.27\text{ARC-C}-0.10\text{HellaSwag}+1.10\text{Winogrande}~{}+\\ &1.98\text{TruthfulQA}-2.82\text{XWinogrande}+2.41\text{HumanEval}-4.93\end{aligned}$
BBH Naive + Greedy $&\phi^{-1}(Y,1.00)\\ =~{}&0.55\log\left(\bar{C}_{\text{Llama-2}}\right)-4.89\\ =~{}&2.65\text{PC1}+3.08\text{PC2}-0.69\text{PC3}-3.22\\ =~{}&1.48\text{MMLU}+1.05\text{ARC-C}+0.69\text{HellaSwag}+0.37\text{Winogrande}~{}+\\ &0.16\text{TruthfulQA}+0.52\text{XWinogrande}+3.58\text{HumanEval}-5.42\end{aligned}$
BBH CoT + Greedy $&\phi^{-1}(Y,1.00)\\ =~{}&0.70\log\left(\bar{C}_{\text{Llama-2}}\right)-5.05\\ =~{}&4.05\text{PC1}+3.86\text{PC2}-2.60\text{PC3}-3.10\\ =~{}&0.99\text{MMLU}+1.29\text{ARC-C}+1.42\text{HellaSwag}+0.42\text{Winogrande}~{}-\\ &0.32\text{TruthfulQA}+1.75\text{XWinogrande}+5.48\text{HumanEval}-6.44\end{aligned}$

### v2

@@ 1 Introduction
Although the high costs of compute scaling laws are unavoidable when optimizing pre-training hyperparameters (e.g., Hoffmann et al. [35]), this is not true of all scaling analyses. In this work, we show that many other types of scaling analyses, such as understanding complex model capabilities (e.g., agentic or “emergent” behaviors) and post-training interventions, can be done using a lower-cost, higher-resolution, and broader-coverage alternative to the standard approach of training (or using) a single family of LMs across compute scales.
Figure 1: Observational scaling laws generalize existing compute scaling laws which directly relate training compute to downstream capabilities (dashed line) by hypothesizing the existence of a low-rank space of LM capabilities that have a log-linear relationship with compute (center), and can be extracted directly from standardized LM benchmarks (left). This enables us to get low-cost, high-resolution scaling predictions of LMs’ complex downstream capabilities from their observable standard benchmark metrics using nearly 80 publicly accessible LMs (left to right).
@@
Through an analysis of existing standardized LM benchmarks (e.g., Open LLM Leaderboard [9]), we find a few such capability measures that have scaling law relationships with compute within model families ($R^{2}>0.9$) (Fig. 3), and with downstream metrics across model families. We call such scaling relationships observational scaling laws as they enable the predictions of complex downstream capabilities from simple observable quantities that we expect to scale with compute (like standardized benchmark performance).
The ability to build scaling laws across a large number of existing LMs from their standard benchmark metrics has significant advantages in cost, resolution, and coverage: Observational scaling incurs no training cost, while leveraging models spanning a much larger compute range than any single model family. It also significantly increases the resolution of scaling laws by virtue of using more models, which is useful for studying nearly discontinuous phenomena like “emergent” capabilities. Finally, observational scaling can combine model families from heterogeneous sources with very different scaling properties (e.g., LLaMA [83] vs StarCoder [46]) which allows us to study how different scaling strategies impact downstream performance and algorithmic interventions.
@@
Emergent capabilities (Sec. 4.1)  There has been an active debate about whether LMs have “emergent” capabilities that discontinuously appear at certain compute thresholds and whether these capabilities can be predicted using small models [90, 74, 56, 36]. The high resolution of observational scaling laws shows that some of these phenomena follow a smooth sigmoid, and can be predicted accurately using small, sub Llama-2 7B models.
Agentic capabilities (Sec. 4.2)  We show that the more high-level, complex capabilities of LMs as agents, as measured by AgentBench [54] and AgentBoard [57], can be predicted with simple benchmark metrics. Our scaling law precisely predicts the performance of GPT-4 using only weaker models (sub GPT-3.5) and identifies programming capabilities as driving agent performance.
Post-training interventions (Sec. 4.3)  We show that our scaling laws can reliably predict the gains of post-training techniques, such as Chain-of-Thought [91] and Self-Consistency [89] at scale, even when we fit our scaling laws on weak models (sub Llama-2 7B).
@@ 2 Related Work
Scaling laws have been generalized beyond pretraining loss to analyze transfer learning [33, 78, 1] and downstream performance [32, 29, 15] across various domains, see Villalobos [86] for a comprehensive review. In particular, there has been evidence suggesting that the few-shot performance of LMs on downstream benchmarks is closely tied to compute measures like model size [13], but whether this is predictable with scaling laws remains debated. Extensive research has explored the difficulties of predicting benchmark performance due to their appearing rapid “emergence” [90, 76, 26], while recent works argued the discontinuity is due to the metrics used [74, 56] or the lack of data points [36] (see Anwar et al. [5] for a survey on this topic). Finnveden [24] and Owen [63] have investigated the use of linear and sigmoidal scaling laws, derived from pretraining loss or computational measures, to extrapolate the benchmark performance. Notably, Owen [63] also utilized publicly available LMs from different families to fit their compute scaling laws despite the potential discrepancies in their compute efficiencies. Recent studies have also more extensively investigated the correlations between the pretraining loss and downstream performance of LMs [93, 37], aiding in the understanding of downstream scaling [25] and emergent capabilities [23] of LMs. On the theory front, Arora and Goyal [6] derived a theory characterizing how performance on complex skills of LMs can be derived as a composition of base skills. While our work shares similar goals in that we aim to understand the downstream, post-training performance of models, we differ in our approach in that we aim to build practical higher-resolution scaling laws using multiple model families and their observable standard benchmark metrics.
@@ 3.1 Generalizing Compute Scaling Laws
We now show that the PC measures $S$ scale log-linearly with training FLOPs within each model family, and can thus be interpreted as a cross-family generalization of compute $C$.
@@
3.4 Fitting Observational Scaling Laws (Eq. 3)
@@
In the multi-model family case, variations in compute efficiency mean that FLOPs and capabilities are no longer log-linear across model families. However, we can map all of the models to a shared, FLOPs-based capability measure of a specific family $f$. The core idea is to represent each model’s capabilities by the following hypothetical: “how many FLOPs ($\bar{C}_{m,f}$) would it take for a model in a family $f$ to match a model $m$”. We call $\bar{C}_{m,f}$ the $f$-equivalent FLOPs for model $m$, as it represents the performance of model $m$ relative to models in the reference model family $f$. This measure can be computed fairly easily as
@@ 4 Validating Observational Scaling Laws
As baselines, we compare our scaling predictions to using existing compute-based scale measures like training FLOPs and model size. We used the mean squared error (MSE) on the holdout set as our main evaluation measure, as the target range is always normalized (0 to 1), and estimating the marginal variance in $R^{2}$ can add additional noise when the test set sizes are small.
@@ 4.1 Predictability of “Emergent” Capabilities
Figure 4: “Emergent” capabilities of LMs can be accurately predicted from weaker models to stronger ones with observational scaling laws, and using PC measures as the predictor provides much more accurate predictions than using compute scale measures like training FLOPs and model size (see Fig. C.10). Two non-arithmetic and two arithmetic tasks from BigBench [75], which are identified as “emergent” in [90], are used for illustration.
@@ 5 Selecting Low-Cost Model Subsets for Practical Scaling Analyses
8 models Llama-2 {7B, 13B, 70B}, Mixtral {8x7B}, Phi {1.5B, 2}, MPT {7B, 30B}
@@
20 models Llama-2 {7B, 13B, 70B}, Mixtral {8x7B}, Qwen {7B, 14B, 72B},
DeepSeek-Coder {1.3B, 6.7B, 33B}, CodeLlama {7B, 13B, 34B, 70B},
MPT {7B, 30B}, Falcon {1B, 7B, 40B, 180B}
@@ 6 Discussion and Other Applications of Observational Scaling
Many existing benchmarks suffer from a limited dynamic range: they either saturate quickly for large models (e.g., HellaSwag, Winogrande) or have completely random performance for small models (e.g., MMLU, GSM8K), see Fig. C.4 for the behavior of each benchmark. In contrast, we find that PC-1 is a smooth capability measure that can be used to compare LMs across many (at least 5) orders of magnitude. This allows us to compare models from heterogeneous sources and of extremely different capabilities on a single, unified scale (Fig. 9). We believe that the high dynamic range of PC1 may make it suitable as an optimization target for pretraining, where architecture or data interventions can be benchmarked against PC-1 at small scales and validated at large scales.
@@ Acknowledgements
Major Changelog
07/02/2024
  • Added clarifications that emphasize the predictions of observational scaling laws are based on standard benchmark metrics instead of training FLOPs.
  • Updated plots to use x-axis with a log10 scale instead of a ln scale for better readability.
  • Updated results with minor numerical deviations after fixing a minor issue.
@@ C.2 Properties of PC measures
In Fig. 9, we have shown that PC-1 can serve as a smooth capability measure for LMs that provide meaningful readouts across many orders of scales (about 5 orders of magnitude). In Fig. C.4, we show that using a single benchmark metric as LM capability measures amy suffer from a limited dynamic range. In particular, they may either saturate quickly for large models (e.g., HellaSwag, Winogrande) or provide random readouts for weak models (e.g., MMLU, GSM8K).
@@ C.3 Robustness Checks
Figure C.8: Comparing different scale measures under different holdout cutoffs on post-training analysis tasks in Sec. 4.3. The training/test data size is varied by changing the FLOPs cutoff and the area under the test error curves (AUE) is used to measure the overall prediction errors. PC measures (with # = 2 or 3) consistently lead to an earlier transition to low prediction error region and much lower AUE compared to compute measures like training FLOPs and model size.
@@ C.6 Model Subset Selection
(d) 20 models
Figure C.16: Prediction results with different numbers of randomly selected models. The prediction performance is much worse than our selection method, even when 20 models are being selected.
@@ C.7 Fited Functional Forms for Preregistration of Predictions
“Emergent” capabilities (Sec. 4.1) Word Unscramble $&\phi^{-1}(Y,1.00)\\ =~{}&2.00\log_{10}(\bar{C}_{\text{Llama-2}})-6.11\\ =~{}&6.74\text{PC1}-3.22\text{PC2}-1.37\text{PC3}-4.93\\ =~{}&1.02\text{MMLU}+3.02\text{ARC-C}+5.73\text{HellaSwag}+2.44\text{Winograd}~{}-\\ &1.06\text{TruthfulQA}+1.21\text{GSM8K}+2.48\text{XWinograd}-0.08\text{HumanEval}-12.28\end{aligned}$
Persian QA $&\phi^{-1}(Y,1.00)\\ =~{}&2.32\log_{10}(\bar{C}_{\text{Llama-2}})-8.43\\ =~{}&2.86\text{PC1}+3.18\text{PC2}-0.19\text{PC3}-5.26\\ =~{}&2.08\text{MMLU}+1.06\text{ARC-C}+1.13\text{HellaSwag}+0.53\text{Winograd}~{}+\\ &0.36\text{TruthfulQA}+2.89\text{GSM8K}+0.66\text{XWinograd}+1.55\text{HumanEval}-7.98\end{aligned}$
3-Digit Substraction $&\phi^{-1}(Y,1.00)\\ =~{}&5.50\log_{10}(\bar{C}_{\text{Llama-2}})-8.92\\ =~{}&5.98\text{PC1}+8.74\text{PC2}+39.55\text{PC3}-4.68\\ =~{}&2.17\text{MMLU}+2.32\text{ARC-C}-3.44\text{HellaSwag}-7.96\text{Winograd}~{}+\\ &0.65\text{TruthfulQA}+34.27\text{XWinograd}+20.39\text{HumanEval}-20.99\end{aligned}$
2-Digit Multiplication $&\phi^{-1}(Y,1.00)\\ =~{}&2.22\log_{10}(\bar{C}_{\text{Llama-2}})-4.45\\ =~{}&3.60\text{PC1}+4.24\text{PC2}+8.05\text{PC3}-2.68\\ =~{}&1.62\text{MMLU}+1.95\text{ARC-C}+0.55\text{HellaSwag}-0.63\text{Winograd}~{}+\\ &0.14\text{TruthfulQA}+6.80\text{XWinograd}+6.52\text{HumanEval}-8.00\end{aligned}$
Agentic capabilities (Sec. 4.2) AgentBench $&\phi^{-1}(Y,0.99)\\ =~{}&\log_{10}(\bar{C}_{\text{Llama-2-Chat}})-5.52\\ =~{}&2.32\text{PC1}+0.79\text{PC2}-2.82\text{PC3}-2.96\\ =~{}&2.34\text{MMLU}+0.82\text{ARC-C}+0.32\text{HellaSwag}+0.54\text{Winogrande}~{}+\\ &0.60\text{TruthfulQA}-0.42\text{GSM8K}+2.63\text{HumanEval}-6.37\end{aligned}$
AgentBoard $&\phi^{-1}(Y,0.97)\\ =~{}&0.98\log_{10}(\bar{C}_{\text{Llama-2-Chat}})-6.60\\ =~{}&3.02\text{PC1}+2.60\text{PC2}+1.17\text{PC3}-2.98\\ =~{}&-0.10\text{MMLU}-0.31\text{ARC-C}-0.55\text{HellaSwag}+0.14\text{Winogrande}~{}+\\ &0.56\text{TruthfulQA}+2.28\text{GSM8K}+3.36\text{HumanEval}-5.06\end{aligned}$
@@
Post-training (analysis Sec. 4.3) GSM Naive + Greedy $&\phi^{-1}(Y,1.00)\\ =~{}&1.09\log_{10}(\bar{C}_{\text{Llama-2}})-4.77\\ =~{}&2.69\text{PC1}+1.55\text{PC2}-0.36\text{PC3}-3.57\\ =~{}&1.53\text{MMLU}+1.30\text{ARC-C}+1.22\text{HellaSwag}+0.75\text{Winograd}~{}+\\ &0.16\text{TruthfulQA}+0.13\text{XWinograd}+1.92\text{HumanEval}-5.97\end{aligned}$
GSM CoT + Greedy $&\phi^{-1}(Y,1.00)\\ =~{}&2.04\log_{10}(\bar{C}_{\text{Llama-2}})-5.53\\ =~{}&2.56\text{PC1}+4.64\text{PC2}+4.21\text{PC3}-2.50\\ =~{}&5.03\text{MMLU}+2.04\text{ARC-C}-0.10\text{HellaSwag}+0.96\text{Winograd}~{}+\\ &1.75\text{TruthfulQA}-2.39\text{XWinograd}+2.58\text{HumanEval}-4.77\end{aligned}$
GSM CoT + SC $&\phi^{-1}(Y,1.00)\\ =~{}&2.19\log_{10}(\bar{C}_{\text{Llama-2}})-5.74\\ =~{}&2.73\text{PC1}+4.82\text{PC2}+4.95\text{PC3}-2.49\\ =~{}&5.58\text{MMLU}+2.27\text{ARC-C}-0.08\text{HellaSwag}+1.11\text{Winograd}~{}+\\ &1.97\text{TruthfulQA}-2.78\text{XWinograd}+2.45\text{HumanEval}-4.95\end{aligned}$
BBH Naive + Greedy $&\phi^{-1}(Y,1.00)\\ =~{}&1.26\log_{10}(\bar{C}_{\text{Llama-2}})-4.87\\ =~{}&2.70\text{PC1}+3.06\text{PC2}-0.84\text{PC3}-3.23\\ =~{}&1.41\text{MMLU}+1.05\text{ARC-C}+0.75\text{HellaSwag}+0.36\text{Winograd}~{}+\\ &0.11\text{TruthfulQA}+0.61\text{XWinograd}+3.63\text{HumanEval}-5.47\end{aligned}$
BBH CoT + Greedy $&\phi^{-1}(Y,1.00)\\ =~{}&1.62\log_{10}(\bar{C}_{\text{Llama-2}})-5.03\\ =~{}&4.20\text{PC1}+3.81\text{PC2}-2.92\text{PC3}-3.12\\ =~{}&0.84\text{MMLU}+1.30\text{ARC-C}+1.57\text{HellaSwag}+0.42\text{Winograd}~{}-\\ &0.44\text{TruthfulQA}+1.96\text{XWinograd}+5.62\text{HumanEval}-6.61\end{aligned}$